OpenStudy (anonymous):

Evaluate the expression. 3 + x - 9 · 6 for x = 2 is -49?

OpenStudy (angela210793):

3+2-54=5-54=-49

OpenStudy (anonymous):

You are correct. It is -49.

OpenStudy (anonymous):

thank you